Q 3.38 For many years chloroform (CHCl3) was used as an inhalation anesthetic in spite of the fact that it is also a toxic substance that may cause severe liver, kidney, and heart damage. Calculate the percent composition by mass of this compound


1
Expert's answer
2021-10-22T03:57:47-0400

The percent by mass of an element in a compound is expressed as

Percent by mass of an element

"=\frac{n \times atomic \; mass \; of \; an \; element}{molecular \;or\; formula\; mass \;of \; compound} \times 100"

Where nis the number of atoms of the element in a molecule or formula unit of the compound.

Use this equation to determine the percent by mass contributed by each element in the compound.

Chloroform (CHCl3) contains one C atom, one H atom and three Cl atoms with the atomic masses 12.01 amu, 1.008 amu and 35.453 amu respectively. The molecular

(12.01 amu) + (1.008 amu) + 3(35.453 amu) = 119.38 amu

For each element, multiply the number of atomic mass, divide by the molecular mass, and multiply by 100 percent.

"\%C = \frac{12.01}{119.38} \times 100 = 10.06\; \% \;C \\

\%H = \frac{1.008}{119.38} \times 100 = 0.844\; \% \;H \\

\%Cl = \frac{3 \times 35.453}{119.38} \times 100 = 89.09 \; \% \;Cl"